If two electrons are each 1.50 × 10-10 m from a proton, find the magnitude and direction of the net electric force they will exert on the proton


1
Expert's answer
2022-01-26T07:41:29-0500
"F=\dfrac 1{4\pi\epsilon_0}\dfrac{2e^2}{r^2}=3.08·10^{-18}\text{ N}."


This force is attractive.